# FedPeWS: Personalized Warmup via Subnetworks for Enhanced Heterogeneous Federated Learning

### v1

@@ Abstract
Statistical data heterogeneity is a significant barrier to convergence in federated learning (FL). While prior work has advanced heterogeneous FL through better optimization objectives, these methods fall short when there is extreme data heterogeneity among collaborating participants. We hypothesize that convergence under extreme data heterogeneity is primarily hindered due to the aggregation of conflicting updates from the participants in the initial collaboration rounds. To overcome this problem, we propose a warmup phase where each participant learns a personalized mask and updates only a subnetwork of the full model. This personalized warmup allows the participants to focus initially on learning specific subnetworks tailored to the heterogeneity of their data. After the warmup phase, the participants revert to standard federated optimization, where all parameters are communicated. We empirically demonstrate that the proposed personalized warmup via subnetworks (FedPeWS) approach improves accuracy and convergence speed over standard federated optimization methods.
@@ 1 Introduction
Consider the scenario where multiple hospitals collaborate to learn a medical image classification model that works across imaging modalities and organs, where the data from each hospital pertains to a different modality (e.g., histopathology, CT, X-ray, ultrasound, etc.) and/or organ (e.g., brain, kidney, colon, etc.). Most of the existing heterogeneous FL algorithms fail when there is such extreme data heterogeneity among collaborating participants, especially when the model is learned from scratch (with random initialization). The main reason for this failure is the high degree of conflicts between the local updates during the initial collaboration rounds. While enforcing a strong regularization constraint on the local updates ([29]) can partially alleviate this problem, it dramatically slows down local learning and hence, convergence speed.
In this work, we explore an alternate approach to minimize the initial conflicts between heterogeneous participants by allowing participants in FL to initially train a partial subnetwork using only their local datasets. This warmup phase enables the participants to focus first on learning their local data well before engaging in broader collaboration. Thus, our proposed approach can be summarized as follows (see Figure 1). Initially, each participant uses a personalized binary mask tailored to their data distributions, allowing them to first learn their local data distributions and optimize their local (sparse) models. During this warmup phase, participants transmit only their masked updates to the server, and this process continues for a certain number of collaboration rounds. At the end of the warmup phase, the participants switch to standard federated optimization methods for subsequent collaboration. Our contributions are as follows:
@@
    We empirically demonstrate the efficacy of the FedPeWS approach under both extreme non-i.i.d. and i.i.d. data scenarios using three datasets: a custom synthetic dataset, a combination of MNIST and CIFAR-10 datasets, and a combination of three distinct medical datasets (PathMNIST, OCTMNIST and TissueMNIST).
@@ 2 Related Work
FL is a distributed learning paradigm that addresses data confidentiality concerns [19], particularly in environments where data can not be centralized due to regulatory or practical reasons [3]. One of the seminal FL algorithms, FedAvg [33], involves participants training models locally on their data and periodically transmitting their model parameters to a central server. The server averages these parameters to update the global model, which is then redistributed to the participants for further local refinement. FedAvg has inspired a plethora of variants and extensions aimed at enhancing performance [20, 29, 35], scalability [12, 1], communication efficiency [43, 38, 15], privacy/confidentiality [41, 8, 43], robustness [26], and fairness [47, 17, 40]. For example, strategies such as weighted averaging or adaptive aggregation have been proposed to accommodate the non-i.i.d. nature of distributed data sources -- a scenario where data is not identically distributed across all participants, which can significantly hinder model performance [29, 45, 20, 31, 44, 7]. Specifically, FedProx [29] addresses data heterogeneity by integrating a proximal term into the FedAvg framework. There is also a body of work that focuses on addressing the heterogeneity problem through personalization-based approaches, utilizing local-centric objectives [11, 13, 49, 39, 6, 28].
@@
Another relevant idea is the Lottery Ticket Hypothesis (LTH) [10], which attempts to identify subnetworks within a larger network. LTH is a model personalization technique, which focuses on sparsifying the network to create a smaller-scale version that improves per-round communication efficiency. In contrast to LTH, our method is directed towards training a shared global model and simultaneously improving convergence speed (reducing number of communication rounds). After LTH, there has been a growing interest in finding sparse and trainable networks at initialization [34, 16, 24]. Recently, in [15], sparse networks were found inside the main model to increase communication efficiency in FL. The proposed FedPM method focuses on finding a subnetwork by freezing the model weights and training for masks on a weight level, in contrast to IST, which works on a neuron level. FedPM utilizes the sigmoid function to obtain probability values from unbounded mask scores and then uses Bernoulli sampling to obtain binary masks. We use a similar approach in our FedPeWS algorithm to learn the neuron-level personalized masks.
@@ 3 Preliminaries
where the components $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ are distributed among $N$ local participants and are expressed in a stochastic format as $f_{i}(x)\coloneqq\mathbb{E}_{\xi\sim\mathcal{D}_{i}}\big{[}F_{i}(x,\xi)\big{]}$. Here, $\mathcal{D}_{i}$ represents the distribution of $\xi$ at participant $i\in[N]\coloneqq\{1,\ldots,N\}$. This problem encapsulates standard empirical risk minimization as a particular case when each $\mathcal{D}_{i}$ is represented by a finite set of $n_{i}$ elements, i.e., $\xi_{i}=\{\xi_{i}^{1},\ldots,\xi_{i}^{n_{i}}\}$. In such cases, $f_{i}$ simplifies to $f_{i}(x,\xi_{i})=\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}F_{i}(x,\xi_{i}^{j})$. Our approach does not impose restrictive assumptions on the data distribution $\mathcal{D}_{i}$. In fact, we specifically focus on the extreme heterogeneous (non-i.i.d.) setting, where $\mathcal{D}_{i}\neq\mathcal{D}_{i^{{}^{\prime}}},\forall~{}i\neq i^{{}^{\prime}}$ and the local optimal solution $x_{i}^{\star}\leftarrow\operatorname*{arg\,min}_{x\in\mathbb{R}^{d}}f_{i}(x)$ might significantly differ from the global minimizer of the objective function in Equation 1.
We are especially interested in the supervised classification task and let $\mathcal{M}_{x}:\mathcal{Z}\rightarrow\mathcal{Y}$ be a classifier parameterized by $x$. Here, $\mathcal{Z}\subseteq\mathbb{R}^{D}$ and $\mathcal{Y}=\{1,2,\ldots,M\}$ denote the input and label spaces, respectively, $D$ is the input dimensionality, $M$ is the number of classes, and $d$ represents the number of parameters in the model $\mathcal{M}$. We set $F_{i}(x,\xi_{i}^{j})=\mathcal{L}(\mathcal{M}_{x}(\mathbf{z}_{i}^{j}),y_{i}^{j})$, where $\mathcal{L}$ is an appropriate loss function and $\xi_{i}^{j}\coloneq(\mathbf{z}_{i}^{j},y_{i}^{j})$ is a labeled training sample such that $\mathbf{z}_{i}^{j}\in\mathcal{Z}$ and $y_{i}^{j}\in\mathcal{Y}$. Furthermore, we mainly focus on the cross-silo FL setting ($N$ is small).
@@
$x_{i}^{t,k}=x_{i}^{t,k-1}-\eta_{\ell}\nabla f_{i}\big{(}x_{i}^{t,k-1},\xi_{i}^{t,k-1}\big{)},\quad x_{i}^{t}=x_{i}^{t,K},\text{ and }x_{g}^{t}=x_{g}^{t-1}-\eta_{g}(x_{g}^{t-1}-\frac{1}{N}\sum_{i=1}^{N}x_{i}^{t}),$ (2)
@@
13:              Optimize over $s$: $\mathcal{L}_{s}=f_{i}\Big{(}x_{i}^{t,k-1}\odot\mathcal{G}\left(s_{i}^{t,k-1}\right),\xi_{i}^{t,k-1}\Big{)}-\lambda\|\sigma\left(s_{i}^{t,k-1}\right)-{\theta_{g\backslash\{i\}}^{t-1}}\|_{2}^{2}$
@@ 4 Proposed FedPeWS Method
Identification of personalized subnetworks: It is not straightforward to directly optimize for the personalized binary (discrete) masks $m_{i}$ in Equation 3. Hence, we make the following design choices. Firstly, personalized masks are learned at the neuron-level and then expanded to the parameter-level. Following IST [50], masks are specifically applied only to the hidden layer neurons, while the head and tail neurons remain unaffected. However, unlike IST, the neuron-level masks are not randomly selected in each collaboration round. Instead, we learn real-valued personalized neuron-level mask score vectors $s_{i}\in\mathbb{R}^{h}$, which in turn can be used to generate the binary masks. Here, $h$ denotes the number of hidden neurons in the classifier $\mathcal{M}$ and $h\ll d$. A higher value of element $s_{i}(\ell)$, $\ell\in[h]$, indicates that the $\ell^{\textrm{th}}$ neuron is more likely to be selected by participant $i$. Let $\mathcal{G}:\mathbb{R}^{h}\rightarrow\{0,1\}^{d}$ be the mask generation function that generates the binary parameter-level masks $m_{i}$ from neuron-level mask score vectors $s_{i}$, i.e., $m_{i}=\mathcal{G}(s_{i})$. $\mathcal{G}$ consists of three steps. Firstly, we convert $s_{i}$ into probabilities by applying a sigmoid function, i.e., $\theta_{i}=\sigma(s_{i})$, where $\theta_{i}\in[0,1]^{h}$ is the mask probability vector and $\sigma$ is the sigmoid function. Next, binary neuron masks $\tilde{m}_{i}$ are obtained by sampling from a Bernoulli distribution with parameter $\theta_{i}$, i.e., $\tilde{m}_{i}(\ell)\sim Bernoulli(\theta_{i}(\ell))$, $\forall~{}\ell\in[h]$. Finally, these binary neuron masks can be directly mapped to the binary parameter-level mask $m_{i}$, i.e., if a neuron is selected, all the weights associated with the selected neuron are also selected. Thus, Equation 3 can be reparameterized as:
@@
$\mathcal{L}_{s}=f_{i}\Big{(}x_{i}^{t,k}\odot\mathcal{G}\left(s_{i}^{t,k}\right),\xi_{i}^{t,k-1}\Big{)}-\lambda\|\sigma\left(s_{i}^{t,k}\right)-\theta_{g\backslash\{i\}}^{t}\|_{2}^{2};\qquad s_{i}^{t,k+1}\leftarrow s_{i}^{t,k}-\eta_{s}\nabla_{s}\mathcal{L}_{s},$ (6)
where $\nabla_{s}$ indicates that the gradient is w.r.t. mask score vector $s$, $\eta_{s}$ is the local learning rate for updating $s$, $\theta_{g}^{t}$ is the global mask probability at round $t$, $\theta_{g\backslash\{i\}}^{t}$ is the global mask probability excluding the probability mask of the current participant $i$, and $\lambda$ is the weight assigned to the mask diversity measure (second term). It is important to note that the personalized masks may not cover all neurons in the network. Maximizing the mask diversity measure encourages personalized masks to deviate as much as possible from the global mask, which facilitates better coverage of all the neurons in the global model. The diversity measure has an upper bound due to the sigmoid function:
$\|\sigma\left(s_{i}^{t,k}\right)-\theta_{g\backslash\{i\}}^{t}\|_{2}^{2}\leq h.$ (7)
Given the difficulty in calculating $\nabla_{s}\mathcal{L}_{s}$ directly due to the discrete nature of Bernoulli sampling, we employ the straight-through estimator (STE) [4] to approximate the gradients, which does not compute the gradient of the given function and passes on the incoming gradient as if the function was an identity function.
@@
where $\nabla_{x}$ indicates that the gradient is w.r.t. weights $x$. The FedPeWS algorithm alternates between these two procedures for $W$ rounds, where $W$ is the number of warmup rounds. At this point, the warmup stops and the participants switch to standard training for $(T-W)$ collaboration rounds. This approach ensures that each participant effectively contributes to the FL process while also tailoring the learning to their specific data distributions. The number of warmup rounds $W$ (or the proportion of warmup rounds $\tau=\frac{W}{T}$) is a key hyperparameter of the FedPeWS algorithm, along with the weight $\lambda$ assigned to the mask diversity loss. While it would be ideal to have a principled method to select these hyperparameters, we use a grid search to tune them, which is currently a limitation.
Use of fixed subnetworks: When the number of participants is small and the data distributions of the participants are known apriori, the server can partition the full model into subnetworks of the same depth and assign a fixed subnetwork to each participant, i.e., $m_{i}^{t}=m_{i}$, $\forall~{}t\in[W]$. Participants transmit only the masked updates back to the server during warmup, which then aggregates these masked parameters and redistributes them in their masked form. For the sake of utility, the server can design personalized masks such that the union of these masks covers all the neurons. This variant of FedPeWS is referred to as FedPeWS-Fixed and follows the same algorithm in Algorithm 1, except for the omission of the highlighted (green) steps.

### v2

@@ Abstract
Statistical data heterogeneity is a significant barrier to convergence in federated learning (FL). While prior work has advanced heterogeneous FL through better optimization objectives, these methods fall short when there is extreme data heterogeneity among collaborating participants. We hypothesize that convergence under extreme data heterogeneity is primarily hindered due to the aggregation of conflicting updates from the participants in the initial collaboration rounds. To overcome this problem, we propose a warmup phase where each participant learns a personalized mask and updates only a subnetwork of the full model. This personalized warmup allows the participants to focus initially on learning specific subnetworks tailored to the heterogeneity of their data. After the warmup phase, the participants revert to standard federated optimization, where all parameters are communicated. We empirically demonstrate that the proposed personalized warmup via subnetworks (FedPeWS) approach improves accuracy and convergence speed over standard federated optimization methods. The code can be found at https://github.com/tnurbek/fedpews.
@@ 1 Introduction
Consider the scenario where multiple hospitals collaborate to learn a medical image classification model that works across imaging modalities and organs, where the data from each hospital pertains to a different modality (e.g., histopathology, CT, X-ray, ultrasound, etc.) and/or organ (e.g., brain, kidney, colon, etc.). Most of the existing heterogeneous FL algorithms fail when there is such extreme data heterogeneity among collaborating participants. The primary reason for this failure is the high degree of conflicts between the local updates and the inherent bias the algorithms introduce, often favoring one modality (dataset) over the others. While enforcing a strong regularization constraint on the local updates [6] can partially alleviate this problem, it dramatically slows down local learning and hence, convergence speed.
In this work, we explore an alternate approach to minimize the initial conflicts between heterogeneous participants by allowing participants in FL to initially train a partial subnetwork using only their local datasets. This warmup phase enables the participants to focus first on learning their local data well before engaging in broader collaboration. Thus, our proposed approach can be summarized as follows (see Figure 1). Initially, each participant uses a personalized binary mask, allowing them to first learn their local data distributions and optimize their local (sparse) models. During this warmup phase, participants transmit only their masked updates to the server, and this process continues for a certain number of collaboration rounds. At the end of the warmup phase, the participants switch to standard federated optimization for subsequent collaboration. Our contributions are as follows:
@@
    We empirically demonstrate the efficacy of the FedPeWS approach under both extreme non-i.i.d. and i.i.d. data scenarios using three datasets: a custom synthetic dataset, a combination of MNIST and CIFAR-10 datasets, a combination of three distinct medical datasets (PathMNIST, OCTMNIST and TissueMNIST), and CIFAR-100 dataset.
@@ 2 Related Work
FL is a distributed learning paradigm that addresses data confidentiality concerns [2], particularly in environments where data can not be centralized due to regulatory or practical reasons [14]. One of the seminal FL algorithms, FedAvg [1], involves participants training models locally on their data and periodically transmitting their model parameters to a central server. The server averages these parameters to update the global model, which is then redistributed to the participants for further local refinement. FedAvg has inspired a plethora of variants and extensions aimed at enhancing performance [7, 6, 15], scalability [16, 17], communication efficiency [18, 19, 20], privacy/confidentiality [21, 22, 18], robustness [23], and fairness [24, 25, 26]. For example, strategies such as weighted averaging or adaptive aggregation have been proposed to accommodate the non-i.i.d. nature of distributed data sources – a scenario where data is not identically distributed across all participants, which can significantly hinder model performance [6, 27, 7, 13, 28, 29]. Specifically, FedProx [6] addresses data heterogeneity by integrating a proximal term into the FedAvg framework. There is also a body of work that focuses on addressing the heterogeneity problem through personalization-based approaches, utilizing local-centric objectives [30, 31, 32, 9, 33, 34].
@@
Shared
Global Model
Level of Mask
Personalization
Learnable
Mask
Learnable
Parameters
FedPM [20] parameter-level $\bm{\times}$
IST [35] neuron-level $\bm{\times}$ (random)
LTH [41] $\bm{\times}$ parameter-level $\bm{\times}$ (pruned)
FedWeIT [32] $\bm{\times}$ parameter-level
FjORD [11] parameter-level $\bm{\times}$ (slimmed)
FedPeWS (ours) neuron-level
@@
Another relevant idea is the Lottery Ticket Hypothesis (LTH) [41], which attempts to identify subnetworks within a larger network. LTH is a model personalization technique, which focuses on sparsifying the network to create a smaller-scale version that improves per-round communication efficiency. In contrast to LTH, our method is directed towards training a shared global model and simultaneously improving convergence speed (reducing the number of communication rounds). After LTH, there has been a growing interest in finding sparse and trainable networks at initialization [42, 43, 44]. In [44, 45], the goal is to learn a personalized model for each client with better communication efficiency. While the sparse subnetworks are learned through pruning in [44], dynamic sparse training-based mask search is employed in [45]. The work by Babakniya et al. [46] focuses on model compression, i.e., learning a sparse server model by identifying a sparse subnetwork within the global model learned via FL. To elaborate further, [44] leverages the lottery ticket hypothesis to find lottery ticket networks inside the main network, and only these networks are communicated during training. The subnetworks are located using pruning techniques, with pruning rates assigned by the server based on hard-coded thresholds for accuracy gains and target pruning rates. At the end of the collaboration, clients obtain personalized models. [45] is a personalized FL technique that employs sparse-to-sparse training with the goal of reducing communication, computation, and memory costs. Sparse models are achieved by applying weight magnitude pruning combined with gradient information computed from a randomly sampled batch. Furthermore, [46] is a sparse learning framework primarily focused on model compression. It identifies subnetworks through sensitivity analysis on a randomly generated mask, which is used to identify a good initial mask for each layer. Our approach fundamentally differs from these methods by aiming to train a single global (dense) model while simultaneously improving the convergence speed (reducing the number of communication rounds required). FedPeWS identifies sparse, personalized subnetworks only during the warmup phase of the training to enable quicker/better convergence under extreme non-IID settings.
Recently, in [20], sparse networks were found inside the main model to increase communication efficiency in FL. The proposed FedPM method focuses on finding a subnetwork by freezing the model weights and training for masks on a weight level, in contrast to IST, which works on a neuron level. FedPM utilizes the sigmoid function to obtain probability values from unbounded mask scores and then uses Bernoulli sampling to obtain binary masks. We use a similar approach in our FedPeWS algorithm to learn the neuron-level personalized masks.
Table 1 presents a comprehensive comparison of our proposed method, FedPeWS, against existing literature for addressing data heterogeneity in federated learning. The comparison highlights key attributes of each method, including the level of mask personalization and whether the masks and parameters are learnable.
Mixture of Experts.
Mixture of Experts (MoE) [47, 48] is a technique that trains multiple specialized models (experts) alongside a gating network that dynamically selects which experts to activate for a given input during both training and inference. Although MoE and our proposed method share the concept of leveraging different model components for learning, there is a fundamental distinction. FedPeWS is not designed to train multiple expert models; rather, it learns a single global model while identifying sparse, personalized subnetworks exclusively during the warmup phase of training. This warmup strategy serves as a crucial initialization step to facilitate efficient federated optimization, improving both convergence speed and generalization in heterogeneous FL environments.
@@ 3 Preliminaries
where the components $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ are distributed among $N$ local participants and are expressed in a stochastic format as $f_{i}(x)\coloneqq\mathbb{E}_{\xi\sim\mathcal{D}_{i}}\big{[}F_{i}(x,\xi)\big{]}$. Here, $\mathcal{D}_{i}$ represents the distribution of $\xi$ at participant $i\in[N]\coloneqq\{1,\ldots,N\}$. This problem encapsulates standard empirical risk minimization as a particular case when each $\mathcal{D}_{i}$ is represented by a finite set of $n_{i}$ elements, i.e., $\xi_{i}=\left\{\xi_{i}^{1},\ldots,\xi_{i}^{n_{i}}\right\}$. In such cases, $f_{i}$ simplifies to $f_{i}(x,\xi_{i})=\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}F_{i}(x,\xi_{i}^{j})$. Our approach does not impose restrictive assumptions on the data distribution $\mathcal{D}_{i}$. In fact, we specifically focus on the extreme heterogeneous (non-i.i.d.) setting, where $\mathcal{D}_{i}\neq\mathcal{D}_{i^{{}^{\prime}}},\forall~{}i\neq i^{{}^{\prime}}$ and the local optimal solution $x_{i}^{\star}\leftarrow\operatorname*{arg\,min}_{x\in\mathbb{R}^{d}}f_{i}(x)$ might significantly differ from the global minimizer of the objective function in Equation 1.
We are especially interested in the supervised classification task and let $\mathcal{M}_{x}:\mathcal{Z}\rightarrow\mathcal{Y}$ be a classifier parameterized by $x$. Here, $\mathcal{Z}\subseteq\mathbb{R}^{D}$ and $\mathcal{Y}=\{1,2,\ldots,M\}$ denote the input and label spaces, respectively, $D$ is the input dimensionality, $M$ is the number of classes, and $d$ represents the number of parameters in the model $\mathcal{M}$. We set $F_{i}\left(x,\xi_{i}^{j}\right)=\mathcal{L}\left(\mathcal{M}_{x}\left(\mathbf{z}_{i}^{j}\right),y_{i}^{j}\right)$, where $\mathcal{L}$ is an appropriate loss function and $\xi_{i}^{j}\coloneq\left(\mathbf{z}_{i}^{j},y_{i}^{j}\right)$ is a labeled training sample such that $\mathbf{z}_{i}^{j}\in\mathcal{Z}$ and $y_{i}^{j}\in\mathcal{Y}$. Furthermore, we mainly focus on the cross-silo FL setting ($N$ is small).
@@
$x_{i}^{t,k}=x_{i}^{t,k-1}-\eta_{\ell}\nabla f_{i}\left(x_{i}^{t,k-1},\xi_{i}^{t,k-1}\right),\quad x_{i}^{t}=x_{i}^{t,K},\text{ and }x_{g}^{t}=x_{g}^{t-1}-\eta_{g}\left(x_{g}^{t-1}-\frac{1}{N}\sum_{i=1}^{N}x_{i}^{t}\right),$ (2)
@@
13:              Optimize over $s$: $\mathcal{L}_{s}=f_{i}\Big{(}x_{i}^{t,k-1}\odot\mathcal{G}\left(s_{i}^{t,k-1}\right),\xi_{i}^{t,k-1}\Big{)}-\lambda\left\|\sigma\left(s_{i}^{t,k-1}\right)-{\theta_{g\backslash\{i\}}^{t-1}}\right\|_{2}^{2}$
@@ 4 Proposed FedPeWS Method
Identification of personalized subnetworks:
It is not straightforward to directly optimize for the personalized binary (discrete) masks $m_{i}$ in Equation 3. Hence, we make the following design choices. Firstly, personalized masks are learned at the neuron-level and then expanded to the parameter-level. Following IST [35], masks are specifically applied only to the hidden layer neurons, while the head and tail neurons remain unaffected. However, unlike IST, the neuron-level masks are not randomly selected in each collaboration round. Instead, we learn real-valued personalized neuron-level mask score vectors $s_{i}\in\mathbb{R}^{h}$, which in turn can be used to generate the binary masks. Here, $h$ denotes the number of hidden neurons in the classifier $\mathcal{M}$ and $h\ll d$. A higher value of element $s_{i}(\ell)$, $\ell\in[h]$, indicates that the $\ell^{\textrm{th}}$ neuron is more likely to be selected by participant $i$. Let $\mathcal{G}:\mathbb{R}^{h}\rightarrow\{0,1\}^{d}$ be the mask generation function that generates the binary parameter-level masks $m_{i}$ from neuron-level mask score vectors $s_{i}$, i.e., $m_{i}=\mathcal{G}(s_{i})$. $\mathcal{G}$ consists of three steps. Firstly, we convert $s_{i}$ into probabilities by applying a sigmoid function, i.e., $\theta_{i}=\sigma(s_{i})$, where $\theta_{i}\in[0,1]^{h}$ is the mask probability vector and $\sigma$ is the sigmoid function. Next, binary neuron masks $\tilde{m}_{i}$ are obtained by sampling from a Bernoulli distribution with parameter $\theta_{i}$, i.e., $\tilde{m}_{i}(\ell)\sim Bernoulli(\theta_{i}(\ell))$, $\forall~{}\ell\in[h]$. Finally, these binary neuron masks can be directly mapped to the binary parameter-level mask $m_{i}$, i.e., if a neuron is selected, all the weights associated with the selected neuron are also selected. Thus, Equation 3 can be reparameterized as:
@@
$\mathcal{L}_{s}=f_{i}\Big{(}x_{i}^{t,k}\odot\mathcal{G}\left(s_{i}^{t,k}\right),\xi_{i}^{t,k-1}\Big{)}-\lambda\left\|\sigma\left(s_{i}^{t,k}\right)-\theta_{g\backslash\{i\}}^{t}\right\|_{2}^{2};\qquad s_{i}^{t,k+1}\leftarrow s_{i}^{t,k}-\eta_{s}\nabla_{s}\mathcal{L}_{s},$ (6)
where $\nabla_{s}$ indicates that the gradient is w.r.t. the mask score vector $s$, $\eta_{s}$ is the local learning rate for updating $s$, $\theta_{g}^{t}$ is the global mask probability at round $t$, $\theta_{g\backslash\{i\}}^{t}$ is the global mask probability excluding the probability mask of the current participant $i$, and $\lambda$ is the weight assigned to the mask diversity measure (second term). It is important to note that the personalized masks may not cover all neurons in the network. Maximizing the mask diversity measure encourages personalized masks to deviate as much as possible from the global mask, which facilitates better coverage of all the neurons in the global model. The diversity measure has an upper bound due to the sigmoid function:
$\left\|\sigma\left(s_{i}^{t,k}\right)-\theta_{g\backslash\{i\}}^{t}\right\|_{2}^{2}\leq h.$ (7)
Given the difficulty in calculating $\nabla_{s}\mathcal{L}_{s}$ directly due to the discrete nature of Bernoulli sampling, we employ the straight-through estimator (STE) [49] to approximate the gradients, which does not compute the gradient of the given function and passes on the incoming gradient as if the function were an identity function.
@@
where $\nabla_{x}$ indicates that the gradient is w.r.t. weights $x$. The FedPeWS algorithm alternates between these two procedures for $W$ rounds, where $W$ is the number of warmup rounds. At this point, the warmup stops and the participants switch to standard training for $(T-W)$ collaboration rounds. This approach ensures that each participant effectively contributes to the FL process while also tailoring the learning to their specific data distributions. The number of warmup rounds $W$ (or the proportion $\tau=\frac{W}{T}$) is a key hyperparameter of the FedPeWS algorithm, along with the weight $\lambda$ assigned to the mask diversity loss. While it would be ideal to have a principled method to select these hyperparameters, we use a grid search to tune them, which is currently a limitation.
Use of fixed subnetworks:
When the number of participants is small and the data distributions of the participants are known apriori, the server can partition the full model into subnetworks of the same depth and assign a fixed subnetwork to each participant, i.e., $m_{i}^{t}=m_{i}$, $\forall~{}t\in[W]$. Participants transmit only the masked updates back to the server during warmup, which then aggregates these masked parameters and redistributes them in their masked form. For the sake of utility, the server can design personalized masks such that the union of these masks covers all the neurons. This variant of FedPeWS is referred to as FedPeWS-Fixed and follows the same algorithm in Algorithm 1, except for the omission of the highlighted (in green color) steps.